# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v1

@@ 1 Introduction
Another example is the general class of regularized problems considered in (Liu et al., 2019), given by
@@
where $g$ and $P_{i}$ are nonsmooth, nonconvex and nonnegative regularizers, and $A_{i}\in{\rm I}\!{\rm R}^{n_{i}\times n}$. This difficult class of optimization problems was addressed by Liu et al. (2019) through the use of the Moreau envelope $M_{\lambda_{i}P_{i}}$ of each $P_{i}$ to obtain an approximate problem:
@@
In this paper, we propose a splitting algorithm that generalizes the DYS algorithm to the four-term optimization problem (1.1) that encompasses a broader range of nonconvex and nonsmooth problem classes, including those described above. We prove that our algorithm is globally subsequentially convergent to stationary points for appropriately chosen stepsizes, with a first-order optimality measure converging at a rate of $1/k$. Another major contribution of this work is the derivation of upper bounds for stepsizes for such convergence results. When specialized to $p=0$, our results yield stepsize estimates for the DYS algorithm that are significantly larger than the existing bounds in (Bian and Zhang, 2021). Up to our knowledge, for this special case of $p=0$, our iteration complexity result is also new for the DYS algorithm. During the final stages of preparing this manuscript, we came across a recently released preprint by Dao et al. (2024) that examines a specific case of our problem in which $L_{{}_{\rm{p}}}=0$ in Assumption 1.1(d). Notably, our analysis yields upper bounds for stepsizes that are considerably better than those reported in their work and an iteration complexity result not covered in theirs.
@@ 2 Preliminaries
Let $\phi:{\rm I}\!{\rm R}^{n}\to(-\infty,\infty]$ be an extended-valued function. The domain of $\phi$ is given by the set $\operatorname*{dom}(\phi)=\{x\in{\rm I}\!{\rm R}^{n}:\phi(x)<\infty\}$. We say that $\phi$ is a proper function if $\operatorname*{dom}(\phi)\neq\emptyset$, and that $\phi$ is closed if it is lower semicontinuous.
$\phi$ is said to be a $\sigma_{\phi}$-convex function if $\phi-\frac{\sigma_{\phi}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex for some $\sigma_{\phi}\in{\rm I}\!{\rm R}$. If $\sigma_{\phi}>0$, then $\phi$ is $\sigma_{\phi}$-strongly convex. If $\sigma_{\phi}\leq 0$, we denote $\rho_{\phi}\coloneqq-\sigma_{\phi}$ and call $\phi$ a $\rho_{\phi}$-weakly convex function. In other words, $\phi$ is $\rho_{\phi}$-weakly convex for $\rho_{\phi}\geq 0$ if $\phi+\frac{\rho_{\phi}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.
@@
$\partial\phi(x)\coloneqq\left\{\xi\in{\rm I}\!{\rm R}^{n}:\exists\{(x^{k},\xi^{k})\}\text{ such that }x^{k}\to x,~{}\xi^{k}\in\hat{\partial}\phi(x^{k}),~{}\phi(x^{k})\to\phi(x),~{}\text{and}~{}\xi^{k}\to\xi\right\},$ (2.1)
where
@@
$\partial\phi(x)=\{\xi\in{\rm I}\!{\rm R}^{n}:\phi(y)\geq\phi(x)+\left\langle\xi,y-x\right\rangle,~{}\forall y\in\mathbb{E}\}.$
@@
$\left\{\xi\in{\rm I}\!{\rm R}^{n}:\exists\{(x^{k},\xi^{k})~{}\text{such that}~{}x^{k}\to x,~{}\xi^{k}\in\partial\phi(x^{k}),~{}\phi(x^{k})\to\phi(x),~{}\text{and}~{}\xi^{k}\to\xi\right\}\subseteq\partial\phi(x).$ (2.2)
@@
and the converse holds if $\phi+\frac{1}{2\gamma}{\left\|{\cdot}\right\|}^{2}$ is convex. We also use the notation $T_{\gamma\phi}:{\rm I}\!{\rm R}^{n}\rightrightarrows{\rm I}\!{\rm R}^{n}$ to denote the forward operator associated with $\phi:{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}$ with stepsize $\gamma>0$, which is given by
@@
where $Id$ is the identity map on ${\rm I}\!{\rm R}^{n}$. Given a point $x$, and a subgradient $\xi\in\partial\phi(x)$, for any $\gamma>0$, we define the quadratic approximation of $\phi$ at $x$ as
@@
When $\phi$ is smooth, we simply write $Q_{\gamma\phi}(w;x)$, and $\xi$ is understood to be equal to $\nabla\phi(x)$.
@@
By Lemma 2.1(a), $\phi$ is $\rho_{\phi}$-weakly convex for some $\rho_{\phi}\in[0,L_{\phi}]$ if $\phi$ is $L_{\phi}$-smooth. If $\phi$ is convex and $L_{\phi}$-smooth, we obtain from Lemma 2.1(b) a familiar property that
@@ 3 Proposed algorithm and its convergence
In this section, we first propose our algorithm, and then provide its convergence analysis under various parameter settings.
@@ 3.1 Proposed algorithm
The above algorithm covers several existing algorithms in the literature.
@@
    $y^{k+1}$ $\in{\rm prox}_{\gamma g}\left(\frac{\gamma}{\alpha}(z^{k}-\alpha\nabla h(z^{k}))+\frac{\gamma}{\beta}T_{\beta p}(y^{k})\right)$
    $z^{k+1}$ $=(1-\tau)z^{k}+\tau y^{k+1}.$
@@ 3.2 Stepsize bounds
From the optimality condition of (3.7), we see that if $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$ for some $x\in{\rm prox}_{\alpha f}(z)$ and $\xi\in\partial p(y)$, then
@@
$\subseteq\left(\frac{1}{\alpha}+\frac{1}{\beta}\right)y^{\scriptscriptstyle+}-\frac{1}{\alpha}T_{\alpha(f+h)}(x)-\frac{1}{\beta}T_{\beta p}(y)+\partial g(y^{\scriptscriptstyle+}).$
@@
for all $(y,z)\in\operatorname*{dom}(g)\times{\rm I}\!{\rm R}^{n}$, $x\in{\rm prox}_{\alpha f}(z)$, $\xi\in\partial p(y)$, and $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$.
@@
$=Q_{\alpha f}(y;x)+\left\langle\nabla f(x)+\frac{1}{\alpha}(y-x),{y^{\scriptscriptstyle+}-y}\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}$
@@
$\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)$ $=Q_{\alpha(f+h)}(y;x)+p(y)+\frac{1}{2\gamma}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}+\left\langle\frac{1}{\alpha}\left(y-T_{\alpha(f+h)}(x)\right)+\xi,y^{\scriptscriptstyle+}-y\right\rangle+g(y^{\scriptscriptstyle+}).$ (3.16)
@@
$\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)$ $\leq Q_{\alpha(f+h)}(y;x)+p(y)+\frac{1}{2\gamma}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}+g(y)-\left(\frac{1}{\gamma}-\frac{\rho_{{}_{\rm{g}}}}{2}\right){\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}.$
@@
Suppose that Assumption 1.1 holds. Let $\rho_{{}_{\rm{f}}}\geq 0$ be such that $f+\frac{\rho_{{}_{\rm{f}}}}{2}\|\cdot\|^{2}$ is convex and $L\geq L_{{}_{\rm{f}}}$ be such that $L-\rho_{{}_{\rm{f}}}>0$. Then
$V_{\Lambda,\xi}(y,z,x)\geq Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle\\ +\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}\\ -\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}$ (3.18)
@@
$Q_{\alpha f}(y^{\scriptscriptstyle+};x)$ $\geq\left(f(\hat{x})+\left\langle\nabla f(\hat{x}),x-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{x-\hat{x}}\right\|}^{2}\right)$
$\qquad+\left\langle\nabla f(x),y^{\scriptscriptstyle+}-x\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2}$
$=f(\hat{x})+\left\langle\nabla f(\hat{x})-\nabla f(x),x-y^{\scriptscriptstyle+}\right\rangle+\left\langle\nabla f(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}$
$\qquad-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{x-\hat{x}}\right\|}^{2}+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2},$
@@
$Q_{\alpha f}(y^{\scriptscriptstyle+};x)$ $\geq Q_{\alpha f}(y^{+};\hat{x})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle-\left(\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x-\hat{x}}\right\|}^{2}$
$\qquad+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}.$ (3.19)
@@
$h(x)+\left\langle\nabla h(x),y^{\scriptscriptstyle+}-x\right\rangle$ $\geq\left(h(\hat{x})-\left\langle\nabla h(x),\hat{x}-x\right\rangle-\frac{L_{{}_{\rm{h}}}}{2}{\left\|{\hat{x}-x}\right\|}^{2}\right)+\left\langle\nabla h(x),y^{\scriptscriptstyle+}-x\right\rangle$
$=\left(h(\hat{x})+\left\langle\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle\right)+\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle-\frac{L_{{}_{\rm{h}}}}{2}{\left\|{\hat{x}-x}\right\|}^{2}$ (3.20)
@@
Regarding the requirement of $\rho_{{}_{\rm{f}}}$ in Lemma 3.4, by Assumption 1.1(a) and Eq. 2.5, we see that there indeed exists $\rho_{{}_{\rm{f}}}\geq 0$ such that $f+\frac{\rho_{{}_{\rm{f}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.
@@
Suppose that Assumption 1.1 holds and $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$. If $\{(x^{k},y^{k},z^{k})\}$ is generated by Algorithm 1 with $\tau\in(0,1]$ and stepsize $\alpha\in(0,\bar{\alpha}]$,
@@
$c(\alpha)\coloneqq\begin{cases}2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)&\text{if}~{}(2-\tau)L_{{}_{\rm{f}}}-2\rho_{{}_{\rm{f}}}\geq\tau L_{{}_{\rm{h}}},\\ 2L_{{}_{\rm{f}}}L_{{}_{\rm{h}}}\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}+\frac{\rho_{{}_{\rm{f}}}\tau(\eta^{*}+\rho_{{}_{\rm{f}}})}{\eta^{*}}\right)\alpha-(2-\tau)&\text{otherwise}.\end{cases}$ (3.27)
@@
$V_{k-1}\geq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})+g(y^{k})+\left\langle\nabla f(x^{k})-\nabla f(x^{k-1})-\frac{1}{\alpha}(x^{k}-x^{k-1}),x^{k-1}-y^{k}\right\rangle\\ +\left\langle\nabla h(x^{k-1})-\nabla h(x^{k}),y^{k}-x^{k}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}\\ -\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.$ (3.29)
@@
[…]
@@
$\qquad-\frac{\alpha}{\tau}\left\langle\nabla h(x^{k})-\nabla h(x^{k-1}),\nabla f(x^{k})-\nabla f(x^{k-1})\right\rangle.$ (3.35)
@@
$\qquad-\frac{\alpha}{\tau}{\left\|{\nabla h(x^{k})-\nabla h(x^{k-1})}\right\|}\cdot{\left\|{\nabla f(x^{k})-\nabla f(x^{k-1})}\right\|}$
@@
$\qquad+\left[\left(\frac{1}{\tau}-\frac{1}{2}\right)\frac{1}{\alpha}-\frac{L_{{}_{\rm{h}}}}{2}-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}-\frac{\alpha}{\tau}L_{{}_{\rm{h}}}L_{{}_{\rm{f}}}\right]{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
$\qquad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}.$ (3.37)

### v2

@@ 1 Introduction
Another example is the general class of regularized problems considered in (Liu et al., 2019a), given by
@@
where $g$ and $P_{i}$ are nonsmooth, nonconvex and nonnegative regularizers, and $A_{i}\in{\rm I}\!{\rm R}^{n_{i}\times n}$. This difficult class of optimization problems was addressed by Liu et al. (2019a) through the use of the Moreau envelope $M_{\lambda_{i}P_{i}}$ of each $P_{i}$ to obtain an approximate problem:
@@
In this paper, we propose a splitting algorithm that generalizes the DYS algorithm to the four-term optimization problem (1.1), which encompasses a broader range of nonconvex and nonsmooth problem classes, including those described above. We prove that our algorithm is globally subsequentially convergent to stationary points for appropriately chosen stepsizes, with a first-order optimality measure converging at a rate of $1/k$. Another major contribution of this work is the derivation of upper bounds for stepsizes for such convergence results. When specialized to $p\equiv 0$, our results yield stepsize estimates for the DYS algorithm that are significantly larger than the existing bounds in (Bian and Zhang, 2021). Up to our knowledge, for this special case of $p\equiv 0$, our iteration complexity result is also new for the DYS algorithm. During the final stages of preparing this manuscript, we came across a recent preprint by Dao et al. (2024) that examines a specific case of our problem in which $L_{{}_{\rm{p}}}=0$ in Assumption 1.1(d). Notably, our analysis yields upper bounds for stepsizes that are considerably larger than those reported in their work.
@@ 2 Preliminaries
Let $\phi:{\rm I}\!{\rm R}^{n}\to(-\infty,\infty]$ be an extended-valued function. The domain of $\phi$ is given by the set $\operatorname*{dom}(\phi)=\{x\in{\rm I}\!{\rm R}^{n}:\phi(x)<\infty\}$. We say that $\phi$ is a proper function if $\operatorname*{dom}(\phi)\neq\emptyset$, and that $\phi$ is closed if it is lower semicontinuous. $\phi$ is said to be a $\sigma_{\phi}$-convex function if $\phi-\frac{\sigma_{\phi}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex for some $\sigma_{\phi}\in{\rm I}\!{\rm R}$. If $\sigma_{\phi}>0$, then $\phi$ is $\sigma_{\phi}$-strongly convex. If $\sigma_{\phi}\leq 0$, we denote $\rho_{\phi}\coloneqq-\sigma_{\phi}$ and call $\phi$ a $\rho_{\phi}$-weakly convex function. In other words, $\phi$ is $\rho_{\phi}$-weakly convex for $\rho_{\phi}\geq 0$ if $\phi+\frac{\rho_{\phi}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.
@@
$\partial\phi(x)\coloneqq\left\{\xi\in{\rm I}\!{\rm R}^{n}:\exists\{(x^{k},\xi^{k})\}\text{ such that }x^{k}\xrightarrow{\phi}x,~{}\xi^{k}\in\hat{\partial}\phi(x^{k}),~{}\text{and}~{}\xi^{k}\to\xi\right\},$ (2.1)
where $x^{k}\xrightarrow{\phi}x$ means $x^{k}\to x$ and $\phi(x^{k})\to\phi(x)$, and
@@
$\partial\phi(x)=\{\xi\in{\rm I}\!{\rm R}^{n}:\phi(y)\geq\phi(x)+\left\langle\xi,y-x\right\rangle,~{}\forall y\in{\rm I}\!{\rm R}^{n}\}.$
@@
$\left\{\xi\in{\rm I}\!{\rm R}^{n}:\exists\{(x^{k},\xi^{k})~{}\text{such that}~{}x^{k}\xrightarrow{\phi}x,~{}\xi^{k}\in\partial\phi(x^{k}),~{}\text{and}~{}\xi^{k}\to\xi\right\}\subseteq\partial\phi(x).$ (2.2)
@@
and the converse holds if $\phi+\frac{1}{2\gamma}{\left\|{\cdot}\right\|}^{2}$ is convex. We also use the notation $T_{\gamma\phi}:{\rm I}\!{\rm R}^{n}\rightrightarrows{\rm I}\!{\rm R}^{n}$ to denote
@@
where $Id$ is the identity map on ${\rm I}\!{\rm R}^{n}$. Given a point $x$ and a subgradient $\xi\in\partial\phi(x)$, for any $\gamma>0$, we define
@@
When $\phi$ is smooth, we write $Q_{\gamma\phi}(w;x)$, and $\xi$ is understood to be equal to $\nabla\phi(x)$.
@@
By Lemma 2.1(a), $\phi$ is $\rho_{\phi}$-weakly convex for some $\rho_{\phi}\in[0,L_{\phi}]$ if $\phi$ is $L_{\phi}$-smooth. If $\phi$ is convex and $L_{\phi}$-smooth, we obtain from Lemma 2.1(b) that
@@ 3 Proposed algorithm and its convergence
In this section, we propose our algorithm and provide its convergence analysis under various parameter settings.
@@ 3.1 Proposed algorithm
The above algorithm covers several algorithms in the literature.
@@
    $y^{k+1}\in{\rm prox}_{\gamma g}\left(\frac{\gamma}{\alpha}(z^{k}-\alpha\nabla h(z^{k}))+\frac{\gamma}{\beta}T_{\beta p}(y^{k})\right),\,z^{k+1}=(1-\tau)z^{k}+\tau y^{k+1}.$
@@ 3.2 Stepsize bounds
Given $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$ for some $x\in{\rm prox}_{\alpha f}(z)$ and $\xi\in\partial p(y)$, we have from the optimality condition of (3.7) that
@@
for all $(y,z)\in\operatorname*{dom}(g)\times{\rm I}\!{\rm R}^{n}$, $x\in{\rm prox}_{\alpha f}(z)$, $\xi\in\partial p(y)$, $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$.
@@
$\begin{split}\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)&=Q_{\alpha(f+h)}(y;x)+p(y)+\frac{1}{2\gamma}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\\ &\quad+\left\langle\frac{1}{\alpha}\left(y-T_{\alpha(f+h)}(x)\right)+\xi,y^{\scriptscriptstyle+}-y\right\rangle+g(y^{\scriptscriptstyle+})\end{split}$ (3.16)
@@
$\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)$ $\leq Q_{\alpha(f+h)}(y;x)+p(y)+g(y)+\left(\frac{1}{2\gamma}-\frac{1}{\gamma}+\frac{\rho_{{}_{\rm{g}}}}{2}\right){\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}.$
@@
Suppose that Assumption 1.1 holds. Let $\rho_{{}_{\rm{f}}}\geq 0$ be such that $f$ is $\rho$-weakly convex and $L\geq L_{{}_{\rm{f}}}$ be such that $L-\rho_{{}_{\rm{f}}}>0$. Then
$\begin{split}&V_{\Lambda,\xi}(y,z,x)\\ \geq&Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle\\ &\quad+\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\end{split}$ (3.18)
@@
$Q_{\alpha f}(y^{\scriptscriptstyle+};x)$ $\geq\left(f(\hat{x})+\left\langle\nabla f(\hat{x}),x-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}\right.$
$\qquad\left.-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{x-\hat{x}}\right\|}^{2}\right)+\left\langle\nabla f(x),y^{\scriptscriptstyle+}-x\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2}$
$=f(\hat{x})+\left\langle\nabla f(\hat{x})-\nabla f(x),x-y^{\scriptscriptstyle+}\right\rangle+\left\langle\nabla f(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle$
$\qquad+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{x-\hat{x}}\right\|}^{2}+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2},$
@@
$Q_{\alpha f}(y^{\scriptscriptstyle+};x)$ $\geq Q_{\alpha f}(y^{+};\hat{x})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle$
$\qquad-\left(\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x-\hat{x}}\right\|}^{2}+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}.$ (3.19)
@@
$h(x)+\left\langle\nabla h(x),y^{\scriptscriptstyle+}-x\right\rangle$
$\geq$ $\left(h(\hat{x})-\left\langle\nabla h(x),\hat{x}-x\right\rangle-\frac{L_{{}_{\rm{h}}}}{2}{\left\|{\hat{x}-x}\right\|}^{2}\right)+\left\langle\nabla h(x),y^{\scriptscriptstyle+}-x\right\rangle$
$=$ $h(\hat{x})+\left\langle\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle-\frac{L_{{}_{\rm{h}}}}{2}{\left\|{\hat{x}-x}\right\|}^{2}$ (3.20)
@@
Regarding the requirement of $\rho_{{}_{\rm{f}}}$ in Lemma 3.4, by Assumption 1.1(a) and Remark 2.2, we see that there indeed exists $\rho_{{}_{\rm{f}}}\geq 0$ such that $f+\frac{\rho_{{}_{\rm{f}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.
@@
Suppose Assumption 1.1 holds and $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$. If $\{(x^{k},y^{k},z^{k})\}$ is generated by Algorithm 1 with $\tau\in(0,1]$ and $\alpha\in(0,\bar{\alpha}]$,
@@
$\begin{split}&c(\alpha)\\ \coloneqq&\begin{cases}2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)&\text{if}~{}(2-\tau)L_{{}_{\rm{f}}}-2\rho_{{}_{\rm{f}}}\geq\tau L_{{}_{\rm{h}}},\\ 2L_{{}_{\rm{f}}}L_{{}_{\rm{h}}}\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}+\frac{\rho_{{}_{\rm{f}}}\tau(\eta^{*}+\rho_{{}_{\rm{f}}})}{\eta^{*}}\right)\alpha-(2-\tau)&\text{otherwise}.\end{cases}\end{split}$ (3.27)
@@
$\begin{split}V_{k-1}&\geq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})+g(y^{k})\\ &\quad+\left\langle\nabla f(x^{k})-\nabla f(x^{k-1})-\frac{1}{\alpha}(x^{k}-x^{k-1}),x^{k-1}-y^{k}\right\rangle\\ &\quad+\left\langle\nabla h(x^{k-1})-\nabla h(x^{k}),y^{k}-x^{k}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.\end{split}$ (3.29)
@@
$~{}\left\langle\nabla f(x^{k})-\nabla f(x^{k-1})-\frac{1}{\alpha}(x^{k}-x^{k-1}),x^{k-1}-y^{k}\right\rangle$
${=}$ $~{}-\frac{\alpha}{\tau}{\left\|{\nabla f(x^{k})-\nabla f(x^{k-1})}\right\|}^{2}+\frac{1}{\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$ (3.34)
@@
$\quad-\frac{\alpha}{\tau}\left\langle\nabla h(x^{k})-\nabla h(x^{k-1}),\nabla f(x^{k})-\nabla f(x^{k-1})\right\rangle.$ (3.35)
@@
$\quad-\frac{\alpha}{\tau}{\left\|{\nabla h(x^{k})-\nabla h(x^{k-1})}\right\|}\cdot{\left\|{\nabla f(x^{k})-\nabla f(x^{k-1})}\right\|}$
@@
$\quad+\left[\left(\frac{1}{\tau}-\frac{1}{2}\right)\frac{1}{\alpha}-\frac{L_{{}_{\rm{h}}}}{2}-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}-\frac{\alpha}{\tau}L_{{}_{\rm{h}}}L_{{}_{\rm{f}}}\right]{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
$\quad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}.$ (3.37)